OpenStudy (mathmath333):

Seven different object must be divided among three people .In how many ways can this be done if one or two of them can get no objects ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{Seven different object must be divided among three}\hspace{.33em}\\~\\ & \normalsize \text{people .In how many ways can this be done}\hspace{.33em}\\~\\ & \normalsize \text{if one or two of them can get no objects ?}\hspace{.33em}\\~\\ & a.)\ 15120 \hspace{.33em}\\~\\ & b.)\ 2187 \hspace{.33em}\\~\\ & c.)\ 3003 \hspace{.33em}\\~\\ & d.)\ 792 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

dbca

ganeshie8 (ganeshie8):

Seven different objects. Notice that each object has 3 different states, each object can go to one of : {person1, person2, person3} So total possible ways = 3*3*3*3*3*3*3 = 3^7 = 2187

OpenStudy (mathmath333):

thanks

OpenStudy (mathmath333):

did this sentence changes any thing "if one or two of them can get no objects"

ganeshie8 (ganeshie8):

yes, that means, one or more can get 0 objects

ganeshie8 (ganeshie8):

obviously all 3 of them cannot get 0 objects because you must give away the 7 objects

ganeshie8 (ganeshie8):

(0, 0, 7) represents the situation : 0 objects to person1 0 objects to person2 7 objects to person3

OpenStudy (mathmath333):

ok thnks